# From Evidence to Verdict: An Agent-Based Forensic Framework for AI-Generated Image Detection

### v1

@@ Abstract
The rapid evolution of AI-generated images poses unprecedented challenges to information integrity and media authenticity. Existing detection approaches suffer from fundamental limitations: traditional classifiers lack interpretability and fail to generalize across evolving generative models, while vision-language models (VLMs), despite their promise, remain constrained to single-shot analysis and pixel-level reasoning. To address these challenges, we introduce AIFo (Agent-based Image Forensics), a novel training-free framework that emulates human forensic investigation through multi-agent collaboration. Unlike conventional methods, our framework employs a set of forensic tools, including reverse image search, metadata extraction, pre-trained classifiers, and VLM analysis, coordinated by specialized LLM-based agents that collect, synthesize, and reason over cross-source evidence. When evidence is conflicting or insufficient, a structured multi-agent debate mechanism allows agents to exchange arguments and reach a reliable conclusion. Furthermore, we enhance the framework with a memory-augmented reasoning module that learns from historical cases to improve future detection accuracy. Our comprehensive evaluation spans 6,000 images across both controlled laboratory settings and challenging real-world scenarios, including images from modern generative platforms and diverse online sources. AIFo achieves 97.05% accuracy, substantially outperforming traditional classifiers and state-of-the-art VLMs. These results demonstrate that agent-based procedural reasoning offers a new paradigm for more robust, interpretable, and adaptable AI-generated image detection.
@@ Introduction
In recent years, image generative models such as GLIDE [37], Imagen [45], DALL·E 2 [43], and Stable Diffusion [44], have advanced rapidly. They can synthesize photorealistic images from natural language in seconds [23, 44, 45]. However, the realism of AI-generated images has raised serious societal concerns. Because AI-generated images can now easily fool human observers, malicious actors are increasingly leveraging them to spread disinformation and impersonate individuals [48]. For example, during the 2024 U.S. presidential election cycle, sophisticated deepfakes (i.e., AI-generated images or videos that convincingly fabricate real people or events) have appeared in campaign ads and on social media, which potentially manipulate public opinions and disrupt the voting behaviors [15]. Beyond elections, AI-generated images also introduce broader risks [48, 50], including misinformation and privacy infringement.
In response to these risks, substantial research has been devoted to the detection of AI-generated images. Current methodologies can be generally classified into two main categories: traditional machine learning classifiers and advanced approaches leveraging large vision language models (VLMs).
@@
Traditional machine learning classifiers typically rely on training convolutional neural networks (CNNs) or transformer-based models to distinguish between real and fake images [53, 59, 54, 19, 17]. Early studies reveal that AI-generated images tend to exhibit shared low-level artifacts, allowing detectors trained on labeled images to identify them [53, 46]. For example, DE-FAKE [46] trains a set of classifiers on AI-generated and real images to learn AI-specific artifacts. However, these works often depend on a limited number of training datasets with fake images produced by only a few specific generative models, which makes them prone to generalizability issues: they perform well on seen data but struggle to generalize to unseen images from new generative models [60, 51]. They also often lack explainability: most models act as black boxes, producing binary outputs without offering human-interpretable justifications [33, 60].
@@
More recently, vision language models (VLMs) have shown promise for more generalizable and explainable detection [60, 57, 33, 28]. Due to the large-scale pre-training, VLMs can be transferred to image detection tasks in a zero-shot or few-shot manner [60, 57, 51], without relying on specialized labeled datasets. Beyond identifying pixel-level artifacts like traditional classifiers, VLMs can also apply semantic and world knowledge in the detection process, e.g., a photo of a flying cat is AI-generated because a flying cat cannot exist in the physical world. In addition, they can provide human-interpretable justifications through prompt engineering for explainable detection [57, 28, 51].
Despite these advancements, both traditional and VLM-based approaches suffer from fundamental limitations compared to human forensic experts. First, they rely heavily on pixel-level image features as the main detection evidence. By contrast, human experts can not only employ off-the-shelf classifiers and VLMs for pixel-level analysis, but also actively seek evidence beyond the images themselves, e.g., Exchangeable Image File Format (EXIF) metadata that contains important information about the image or online contextual information to make a comprehensive decision. Second, both approaches treat image detection as a single-shot classification task, using a fixed model or pipeline. Human experts, on the other hand, approach image detection as a dynamic reasoning process: they flexibly employ different tools and iteratively refine their judgments based on the gathered evidence. Finally, human experts can improve over time through accumulated experience, enabling them to adapt to new generative models and real-world scenarios, whereas fixed models remain static unless frequently fine-tuned on new datasets. Motivated by this, we aim to explore a new paradigm that combines the advantages of existing classifiers and VLMs with the capabilities of human experts for AI-generated image detection.
Our Work. Our approach fundamentally differs from conventional methods based on static classifiers or single VLM. Rather than developing another standalone image classifier, we design a cognitive AI agent system that automates the entire human-like forensic workflow, simulating reasoning and decision-making in the detection of AI-generated images. We present AIFo (Agent-based Image Forensics), an LLM-based multi-agent framework that integrates functionalities such as visual recognition, semantic understanding, and provenance analysis. AIFo consists of a forensic Toolbox, an Evidence Gatherer, a Reasoning Agent, and a Debate module. Given a test image, the Evidence Gatherer first applies tools from the Toolbox and aggregates the resulting evidence. The Reasoning Agent then evaluates the quality of the collected evidence, i.e., whether it is sufficient and consistent enough to support a reliable judgment. If so, the Reasoning Agent produces a final decision and explains why the image is classified as AI-generated or real. Otherwise, we introduce a Debate module [21, 35, 36, 34, 49] to handle cases where the evidence is insufficient or conflicting. In this module, two Debate Agents exchange arguments over multiple rounds, each adopting an opposing stance (for or against the claim that the image is AI-generated), while a Judge Agent oversees the debate process and produces the final judgment. Additionally, we demonstrate the potential of our framework to continuously and progressively enhance its detection effectiveness by learning from historical testing data. By incorporating a memory module that stores all testing history, the framework behaves like a human expert, accumulating experience and improving performance over time. Overall, such a cognitive AI agent framework offers a new paradigm for AI-generated image detection.
Main Findings. We evaluate AIFo on a dataset of 6,000 AI-generated and real images, comprising 3,000 samples from five established benchmarks (e.g., Flickr30K [40], GenImage [61], and FakeBench [33]) and 3,000 in-the-wild images collected from six online platforms. AIFo is benchmarked against a range of baselines methods, including traditional classifiers such as CNNSpot [53], DE-FAKE [46], and state-of-the-art VLMs (e.g., GPT-4.1 [8] and GPT-4o [39]). To ensure a fair comparison, we disable the memory module and prevent our framework from learning during the main evaluation. Even under these conditions, AIFo achieves the best overall performance, reaching 0.9705 accuracy and surpassing GPT-4o (0.9483), GPT-4.1 (0.9416), and other baselines. It also maintains higher robustness (0.9047–0.9690) than GPT-4o (0.8818–0.9462) under three types of perturbations. To demonstrate the effectiveness of the memory module, we conduct a case study on 50 images that are misclassified in the main evaluation. The results show that when the memory module stores similar historical cases, approximately 40% of these errors are successfully corrected.
The main contributions of our work are as follows:
  • Human-Like Procedural Reasoning. We present the first agentic framework (AIFo) that simulates human-like procedural reasoning for AI-generated image detection. Our framework combines the strengths of conventional classifiers and VLMs, employing them as complementary tools, and incorporates human-like reasoning processes such as integrating multiple sources of evidence, debating conflicting evidence, and improving through accumulated experience. This highlights a transition from static classification toward a dynamic reasoning process in AI-generated image detection.
  • Comprehensive Evaluation. We contribute a valuable benchmark dataset covering 6000 images from both existing benchmark datasets and in-the-wild images from internet platforms. This dataset serves as the foundation for rigorous evaluation, through which we demonstrate the superior performance, generalizability, and robustness of our AIFo compared to state-of-the-art baselines.
  • Training-Free Core and Cross-Model Generalizability. Unlike conventional detectors that require large-scale training data and frequent updates, the core of AIFo is training-free and designed to generalize across evolving generative models. It leverages generalizable forensic evidence rather than model-specific features, enabling robust detection of AI-generated images. An optional memory module further enhances performance over time by accumulating historical cases.
Threat Model
Detector’s Goals
Distinguishing Between AI-Generated and Real Images. The detector aims to determine whether a given image is AI-generated or a real image. In this work, AI-generated images include those produced by generative models like text-to-image diffusion models, as well as AI-edited images, such as cartoons generated from real photographs and images replaced with AI-generated backgrounds. Real images refer to those captured by physical cameras or manually created/drawn by human artists. Notably, if a real image has been edited or post-processed by humans (without the aid of AI models), e.g., by applying filters, adjusting color contrast, or increasing brightness, we still consider it a real image.
Enhanced Interpretability. Unlike conventional binary detectors, our detector aims not only to provide a decision but also to offer human-interpretable justifications. For example, if our detector identifies an image as an AI-generated image, it should also point out evidence that supports this decision, such as unreasonable object placement (semantic-level evidence), distorted lighting in the image (pixel-level evidence), editing history (metadata-level evidence), AI watermarks (source-level evidence), etc. Providing such explanations helps users understand, verify, and trust the detector’s decisions.
Cross-Model Generalizability. Given the rapid evolution of text-to-image generative models, new architectures and generated images are continually emerging, which may degrade the performance of detectors trained on previous generative models. Therefore, the detector should be robust to various model iterations and new architectures, and capable of leveraging generalizable evidence to assess image provenance, rather than relying solely on model-specific features.
Training-Free Design. Furthermore, traditional AI image detection frameworks typically depend on collecting large-scale training datasets of AI-generated images. As generative models evolve, these frameworks require continual updates to their training data, incurring significant maintenance costs. In contrast, the core of our detector is designed to be training-free: it does not require any dedicated training set, and can be directly applied to detect AI-generated images from any model, including those unseen during development.
Detector’s Capabilities
Our detector is assumed to have access to the following resources and conditions:
  • Image-File Access. The detector can access the image file itself, including its metadata if available (e.g., camera details, timestamps, and editing history).
  • External Forensic Tools. The detector is permitted to utilize a suite of external forensic tools to analyze the image and gather supporting evidence. This includes pretrained AI detection models, reverse image search engines, and vision-language models.
@@
Unlike existing approaches that rely solely on pre-trained classifiers or VLMs, our main design principle is to emulate the procedural reasoning processes employed by human forensic experts [22, 38], while still leveraging the strengths of existing approaches. To this end, we leverage LLMs as autonomous agents to develop an agentic framework that incorporates the key capabilities of human forensic experts. First, the framework is designed to employ and coordinate a suite of forensic tools, such as pre-trained classifiers, metadata extraction, reverse image search, and VLM-based analysis. Second, rather than relying on a single fixed source of evidence, the framework should be able to cross-check evidence from multiple tools before forming the final decision. Next, when the collected evidence indicates conflicting signals, the framework is expected to resolve these conflicts through human-like reasoning activities, such as assessing the reliability of evidence sources and engaging in structured debate among the expert team. The core framework is training-free, building on top of pre-trained classifiers and state-of-the-art VLMs as forensic tools. At the same time, an optional memory module allows the framework to accumulate historical cases and progressively improve performance. Altogether, we aim to build a training-free, interpretable, and robust framework for AI-generated image detection, which not only leverages existing detection approaches but also possesses key capabilities of human forensic experts, e.g., integrating various evidence, evaluating their reliability, and resolving conflicts among them.
@@
Figure 1: High-level overview of our proposed AIFo.
Our proposed framework, Agent-based Image Forensics (AIFo), is a training-free, LLM-based multi-agent system for AI-generated image detection, designed to mimic the procedural reasoning workflow of human forensic experts. As illustrated in Figure 1, the core of our framework [^1] We introduce the optional memory module in Section 5. consists of a Toolbox $\mathcal{T}=\{T_{1},T_{2},\dots,T_{n}\}$ containing $n$ different forensic tools, an Evidence Gatherer Agent $A_{EG}$, a Reasoning Agent $A_{R}$, two Debate Agent $A_{D1}$ and $A_{D2}$ and a Judge Agent $A_{J}$. Given an input image of unknown authenticity from the set of all images, $I\in\mathcal{V}$, the Evidence Gatherer $A_{1}$ first employs tools from the Toolbox. Specifically, the Toolbox includes four main types of forensic tools: (1) Image Reverse Search to locate visually similar images and retrieve their provenance information from online sources; (2) Metadata Extraction to recover embedded EXIF or generated metadata from the image file; (3) VLM-Based Analysis to perform semantic-level reasoning and contextual understanding of the image leveraging VLMs; and (4) Pre-Trained Classifiers to apply pre-trained models for binary classification of the image.
Then, the Evidence Gatherer Agent executes each tool $T_{i}\in\mathcal{T}$ to produce a piece of evidence:
$e_{i}=T_{i}(I),\quad T_{i}:\mathcal{V}\rightarrow\mathcal{E}_{i},$
where $\mathcal{E}_{i}$ denotes the evidence space for tool $T_{i}$. For example, the metadata extraction tool may return detailed EXIF information embedded in the image file, such as "EXIF:Model": "Canon EOS 5D Mark IV" or "EXIF:LensModel": "EF24-70mm f/2.8L II USM", which can serve as important evidence for authenticity assessment. The complete evidence set is:
$E=\{e_{i}\mid T_{i}\in\mathcal{T}\}.$
Next, the Reasoning Agent $A_{R}$ examines the entire evidence set $\mathcal{E}$ and determines whether the evidence is sufficient and consistent enough to make a reliable decision. Evidence sufficiency refers to whether the collected evidence covers enough aspects of the input image, e.g., without missing key metadata or reverse search results. Evidence consistency measures the extent to which different pieces of evidence point toward the same conclusion, e.g., when the majority of evidence supports that the image is AI-generated. If the evidence is considered sufficient and consistent enough, the Reasoning Agent generates a final judgment $D\in\mathcal{D}$ and a human-readable explanation $R\in\mathcal{R}$:
$A_{R}:\mathcal{E}\xrightarrow{}\mathcal{D}\times\mathcal{R}.$
Otherwise, the system initiates a multi-round debate process. Two Debate Agents $A_{D1}$ and $A_{D2}$ take opposing stances (pro vs. contra) regarding whether the image is AI-generated or real. They exchange arguments over $n$ rounds, and a Judge Agent $A_{J}$ observes the debate history $H$ together with the tool-derived evidence $\mathcal{E}$, and produces the final judgment $D\in\mathcal{D}$ along with an explanation $R\in\mathcal{R}$:
$A_{J}:(H,\mathcal{E})\rightarrow\mathcal{D}\times\mathcal{R}.$
This hierarchical reasoning-debate framework ensures that decisions are made either directly from sufficient evidence or through a careful debating process that simulates human forensic reasoning when the evidence is inconclusive or conflicting.
LLM-Based Agents
The LLM-based agents act as the central reasoning and orchestration units of our multi-agent framework, responsible for applying forensic tools, evaluating the reliability of collected evidence, and making decisions based on existing evidence. All agents are instantiated from a LLM (GPT-4.1-2025-04-14), with specialized prompts defining their respective roles and responsibilities. The detailed prompts used for each agent are provided in Appendix A.
The Evidence Gatherer has full access to the Toolbox and is responsible for invoking the forensic tools to collect evidence. Its decision-making process is guided by a carefully engineered prompt that contains three critical components: (1) a clear definition of the agent’s task, which is acting as an AI image forensics expert to determine whether the input image is AI-generated or real; (2) rigorous definitions of “AI-generated image” and “real image” (see Section 2.1) to ensure consistent interpretation during analysis; (3) a comprehensive description of all available tools, including their capabilities. With these explicit task definitions and resource descriptions, the LLM gains both the contextual understanding and the operational awareness to invoke the tools and collect results.
The Reasoning Agent is responsible for conducting an initial assessment of the evidence collected by the evidence gatherer. It first evaluates whether the evidence set is sufficient and consistent enough. This evidence set check is implemented via a dedicated prompt, which instructs the LLM to internally assess the completeness (sufficiency) and consistency of the evidence and return a binary decision (true / false). If true, the reasoning agent proceeds to generate a final judgment and an explanation with another prompt that: (1) defines its responsibility to synthesize the collected evidence and produce a final binary judgment (AI-generated or real) with an accompanying explanation; (2) instructs the agent to evaluate the reliability of each evidence source, taking into account the confidence level, potential biases, and the trustworthiness of the generating tool; (3) emphasizes the importance of generating a human-readable, logically structured reasoning chain. If the evidence is insufficient or conflicting, the reasoning agent abstains from producing a final decision and instead triggers the multi-agent debate mechanism.
The multi-agent debate mechanism is designed to resolve ambiguity and enhance the robustness of the final judgment when the evidence is inconclusive or conflicting.
The Debate Agents are responsible for engaging in a structured multi-round debate when the evidence is deemed insufficient or inconsistent. The debate process lasts up to $n$ rounds. In the first round, the two debate agents are guided by designed prompts to take opposing stances: one agent provides arguments supporting that the image is AI-generated based on collected evidence, while the other provides arguments supporting that the image is real. In subsequent rounds, both agents are instructed to refine or strengthen their reasoning based on the arguments made by the opposing side in the previous round. Through this iterative exchange, the debate agents progressively sharpen their analyses and solve potential conflicts in the evidence interpretation.
A Judge Agent is responsible for supervising the debate process. At the end of each round, the Judge Agent internally evaluates whether the debate has reached sufficient coverage and clarity to make a final decision. It may decide to terminate the debate early if the arguments are deemed sufficient, or allow the debate to proceed for additional rounds if further clarification is needed. Once the debate concludes, the Judge Agent generates the final judgment and an explanation. Unlike the Reasoning Agent, which relies solely on the collected evidence, the Judge Agent is guided to synthesize both the tool-derived evidence and the debate history. This design ensures that the final decision integrates multiple sources of evidence with structured deliberation, leading to a transparent and defensible forensic conclusion.
Forensic Tools

### v2

@@ Abstract
The rapid evolution of AI-generated images poses growing challenges to information integrity and media authenticity. Existing detection approaches face limitations in robustness, interpretability, and generalization across diverse generative models, particularly when relying on a single source of visual evidence. We introduce AIFo (Agent-based Image Forensics), a training-free framework that formulates AI-generated image detection as a multi-stage forensic analysis process through multi-agent collaboration. The framework integrates a set of forensic tools, including reverse image search, metadata extraction, pre-trained classifiers, and vision-language model analysis, and resolves insufficient or conflicting evidence through a structured multi-agent debate mechanism. An optional memory-augmented module further enables the framework to incorporate information from historical cases. We evaluate AIFo on a benchmark of 6,000 images spanning controlled laboratory settings and challenging real-world scenarios, where it achieves 97.05% accuracy and consistently outperforms traditional classifiers and strong vision-language model baselines. These findings demonstrate the effectiveness of agent-based procedural reasoning for AI-generated image detection.
@@ Introduction
Figure 1: High-level overview of our proposed AIFo.
Recent years have witnessed rapid advancements in image generative models, ranging from early works such as GLIDE [NDRSMMSC21], Imagen [SCSLWDGAMLSHFN22], DALL·E 2 [RDNCC22], and Stable Diffusion [RBLEO22] to more recent foundation models including HunyuanImage [CCCCCDDGGGo25] and Qwen-Image [WLZLGYYBXCCTZWYYCLLZMWNCCPQWWYWFXWZZWCL25]. They can synthesize photorealistic images from natural language in seconds. However, the realism of AI-generated images has raised serious societal concerns, as they can be used for disinformation, impersonation, and privacy infringement, undermining public trust [SWAZ23, V20]. In response to these risks, substantial research has been devoted to the detection of AI-generated images. Current methodologies can be generally classified into two main categories: traditional machine learning classifiers and approaches leveraging large vision-language models (VLMs).
@@
Traditional machine learning classifiers typically rely on training convolutional neural networks (CNNs) or transformer-based models to distinguish between real and fake images [WWZOE20, ZXLQZ24, WBZWHCL23, CPNV24, BLYXLHC23]. Early studies reveal that AI-generated images tend to exhibit shared low-level artifacts, allowing detectors trained on labeled images to identify them [WWZOE20, SLYZ22]. For example, DE-FAKE [SLYZ22] trains a set of classifiers on AI-generated and real images to learn AI-specific artifacts. While effective under controlled settings, such approaches often struggle to generalize to unseen generative models and provide limited insight into how individual cues contribute to the final decision [ZLWSJYDSWJ25, WDWJYZZQXZM25, LLWLWRL24].
@@
More recently, VLMs have shown promise for more generalizable and explainable detection [ZLWSJYDSWJ25, YFGFXC25, LLWLWRL24, JHZCLZWZZ25]. Due to the large-scale pre-training, VLMs can be transferred to image detection tasks in a zero-shot or few-shot manner [ZLWSJYDSWJ25, YFGFXC25, WDWJYZZQXZM25], without relying on specialized labeled datasets. Moreover, through prompt engineering, VLMs are able to produce interpretable justifications alongside their predictions [YFGFXC25, JHZCLZWZZ25, WDWJYZZQXZM25].
Despite these advancements, both traditional and VLM-based methods remain limited compared to human forensic experts. Human experts rarely rely on a single visual cue, instead integrating diverse evidence sources and iteratively refining their conclusions as new information emerges. By contrast, most existing methods perform fixed classification and lack explicit mechanisms to assess evidence sufficiency, resolve conflicting signals, or accumulate experience over time. Motivated by this, we seek a new paradigm that integrates the strengths of classifiers and VLMs with the procedural reasoning capabilities exhibited by human experts for AI-generated image detection.
Instead of developing another image classifier, we argue that AI-generated image detection should be viewed as an evidence-driven decision-making problem. Therefore, we adopt a procedural reasoning formulation inspired by human forensic workflows, in which a verdict is reached through multiple stages of reasoning rather than a single forward pass. We present AIFo (Agent-based Image Forensics), a training-free, multi-agent framework that coordinates multiple forensic tools through specialized agent roles. As shown in Figure 1, AIFo structures the detection process into distinct stages, including evidence collection, evidence assessment, and an optional debate stage. An Evidence Gatherer Agent invokes tools from a forensic toolbox and aggregates their outputs, while a Reasoning Agent evaluates whether the collected evidence is sufficient and consistent to support a reliable decision. When evidence is ambiguous or contradictory, a debate mechanism is activated, allowing opposing hypotheses to be examined under the supervision of a Judge Agent. We further introduce a memory module that stores historical cases, enabling the framework to accumulate experience and improve performance over time. We evaluate AIFo on a dataset of 6,000 AI-generated and real images, comprising 3,000 samples from five established benchmarks (e.g., Flickr30K [PWCCHL15], GenImage [ZCYHLLTHHW23], FakeBench [LLWLWRL24], among others) and 3,000 real-world images from six online platforms. Across these settings, AIFo consistently outperforms traditional classifiers such as CNNSpot [WWZOE20] and DE-FAKE [SLYZ22], as well as strong VLM baselines (e.g., GPT series [O23]), achieving higher accuracy and robustness under common perturbations.
The main contributions of our work are: (1) We reformulate AI-generated image detection as an evidence-driven forensic decision problem, highlighting the role of evidence sufficiency, consistency, and conflict resolution under uncertainty. Building on this, we present an agent-based procedural reasoning framework that implements key stages of human forensic analysis, including evidence aggregation, evidence assessment, and structured conflict resolution. (2) Our proposed framework AIFo is a training-free system that integrates conventional classifiers and VLMs as complementary evidence sources rather than standalone detectors, with a modular design that enables robust generalization to evolving generative models. (3) We conduct extensive empirical evaluations on a benchmark of 6,000 images combining public datasets and real-world samples, and demonstrate the superior performance, generalizability, and robustness of AIFo over state-of-the-art baselines.
Related Work
Fake Image Detection. Early approaches to AI-generated image detection mainly rely on training image classifiers using machine learning, often using datasets produced by specific generative models. Over time, research has shifted towards more generalizable and robust detection strategies, including leveraging the visual capabilities of large multimodal models (LMMs) such as CLIP [RKHRGASAMCKS21]. CNNSpot [WWZOE20] proposed one of the first universal detectors independent of generator architecture or dataset. De-Fake [SLYZ22] combined image content and textual prompts using CLIP-based hybrid training. DIRE [WBZWHCL23] and ZeroFake [STLBZ24] exploit intrinsic differences between real and fake images revealed during the diffusion model reconstruction process to build detection models with improved generalization. PatchCraft [ZXLQZ24] focuses on local texture patches, identifying subtle artifacts left by generative models in fine-grained regions. AIDE [YLCHJHX24] detects AI-generated images by selecting highest and lowest frequency patches through a mixture-of-experts architecture.
Recent research leverages VLMs for detection, such as AntifakePrompt [CYCY23], Jia et al. [JLZCYJHLWL24], and Ji et al. [JHZCLZWZZ25]. AIGI-Holmes [ZLWSJYDSWJ25] proposes a complete framework to train a VLM for explainable and generalizable detection, aiming to produce human-verifiable justifications. The work by Yu et al. [YFGFXC25] develops a framework to enhance generalization and explainability by using a knowledge-guided detector and a forgery-aware prompt learner. FakeBench [LLWLWRL24] and DFBench [WDWJYZZQXZM25] introduced a large-scale benchmark to rigorously test the detection performance of LMMs against a wide range of modern generative models.
Despite these advances, existing methods still mainly rely on internal, pixel-level visual features, overlooking complementary external evidence critical to forensic reasoning. In contrast, our agent-based framework emulates human investigative workflows, leverages external tools and reason over different evidence, which allows for a robust and explainable framework for AI-generated image detection.
LLM-Based Multi-Agent Frameworks. Recent advances in LLMs have catalyzed the development of agentic frameworks that simulate complex human-like workflows by coordinating multiple specialized agents. ReAct [YZYDSNC23] introduces a framework that switches between reasoning and acting within LLMs. Other works such as MetaGPT [HZCZCWZWYLZRXWS24] simulate various roles in a software company and build a multi-agent software development framework. Moreover, multi-agent frameworks have also been applied to adversarial defense [ZWZWW24], harmful content detection [LFLWD25], bias identification in generative models [WBYXCMZCW23], fake news verification [LZM24], and misinformation evaluation [HYYX25]. However, existing approaches primarily emphasize predictive accuracy and explainability, while leaving evidence aggregation and conflict resolution largely implicit. We therefore present the first training-free multi-agent framework tailored to this task, combining the strengths of existing detectors and LLM agents for adaptive, explainable AI image forensics.
@@
Instead of treating AI-generated image detection as a single-shot visual classification problem, we formulate it as an evidence-driven reasoning process that operates under heterogeneous, incomplete, and sometimes conflicting evidence, emulating the procedural reasoning of human forensic experts [FH24, NLBFPCL24]. Our framework follows a multi-stage reasoning procedure: (1) collect evidence from multiple sources, (2) assess whether the evidence is sufficient or consistent to justify a decision, (3) resolve conflicts through debate mechanism, and (4) produce a verdict with traceable rationale.
@@
Given an input image $I\in\mathcal{V}$, the objective is to produce a binary verdict $D\in\mathcal{D}=\{\texttt{AI},\texttt{Real}\}$ together with a rationale $R\in\mathcal{R}$. Rather than predicting $D$ directly from pixel-level features, the framework reasons over a set of heterogeneous evidence items $E=\{e_{1},e_{2},\dots,e_{m}\}$, where each $e_{k}$ is an evidence item produced from a distinct evidence channel (e.g., provenance, metadata, model predictions, semantic analysis). Each evidence item is represented as a structured record $e_{k}=\langle t_{k},\;s_{k},\;\xi_{k}\rangle$, where $t_{k}$ denotes the evidence type, $s_{k}\in\{\texttt{AI},\texttt{Real},\texttt{Unknown}\}$ indicates the qualitative indication suggested by the evidence source, and $\xi_{k}$ contains the structured output produced by the corresponding tool. In AIFo, the process of collecting evidence is carried out by an Evidence Gatherer Agent, which invokes a set of forensic tools and normalizes their outputs into the unified evidence representation above.
The complete evidence set $E$ then serves as the input to a reasoning module, whose role is to qualitatively assess whether the evidence provides adequate support for a reliable decision and to identify agreements or contradictions across evidence items. The assessment such as evidence sufficiency and consistency are not computed as predefined metrics, but are instead explicitly reasoned about by a Reasoning Agent based on the available evidence and its contextual descriptions. Evidence sufficiency refers to whether the currently available evidence provides adequate support for a confident decision, and evidence consistency captures the degree to which different evidence sources support compatible conclusions. If the reasoning module concludes that the evidence is insufficient or contains unresolved conflicts, the framework defers commitment and invokes a structured conflict resolution procedure. Inspired by previous work [DLTTM24, LHJWWWYST24, LFLWD25, LDZHGLI24, SGDBP24], the framework adopts a multi-agent debate (MAD) mechanism to explicitly examine competing arguments rather than aggregating evidence. The debate is organized around two opposing hypotheses that the input image is AI-generated or real, and two agents are instantiated with complementary roles, each tasked with constructing the argument in support of one hypothesis using the same set of collected evidence. They exchange arguments over up to $n$ rounds, refining their reasoning based on prior exchanges at each round, and the debate continues until the supervising Judge Agent determines that one hypothesis is more defensible or the maximum number of rounds is reached. At termination, the Judge Agent synthesizes the debate history $H$ together with the original evidence $E$ to produce a final judgment $D\in\mathcal{D}$ and a corresponding explanation $R\in\mathcal{R}$: $A_{J}:(H,E)\rightarrow\mathcal{D}\times\mathcal{R}$. We instantiate the above procedural reasoning framework in AIFo, a training-free, LLM-based multi-agent system that maps each stage of the decision process to a specialized agent role. Each agent operates under a clearly defined role with restricted responsibilities and are guided by structured prompts.
Evidence Providers and Tool Interfaces